Sol Summary – April 18th

Crew 280 Sol Summary Report 18-04-2023

Sol: 2

Summary title: "Hypatia on Air: Mars, Camera, Action"

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: The crew woke up very early in order to welcome a media visit from Earth. Before this visit, some members had the opportunity to take their first shower, which was a nice reminder of the luxuries that we have on Earth and that we sometimes take for granted. After everyone got ready, we had breakfast and started working on our experiments.

At about 11:15AM, Mariona Badenas-Agusti (Crew Commander), Carla Conejo (Executive Officer), Neus Sabaté (Mission Specialist), and Núria Jar (Journalist) conducted an EVA. The original plan was to drive to Cowboy Corner, but we requested a change to the initial plan due to the high winds around the station. Our plan was approved, and the EVA team drove to Marble Ritual instead, using Curiosity and Spirit rovers.

The goal of today’s EVA was to take geological samples for Neus’ research experiments and get more familiar with the EVA protocols and spacesuits. At the same time, the EVA offered the media team a chance to obtain beautiful pictures and high-quality drone videos. We are looking forward to seeing their results.

At around 1PM, the EVA group returned to the Hab and had a delicious meal with the rest of the team. Laia Ribas (Greenhab Officer) and Cesca Cufí (Crew Engineer) prepared pasta with roasted vegetables and parmesan cheese and fruit. We then rested a bit before everyone continued to prepare their scientific experiments.

Soon after 5PM, the media team left the Hab and the crew participated in Carla’s psychological experiment. This evening, we will get in touch with Mission Support, have our second dinner together, and organize a debate about the future of Hypatia and our Association.

Look ahead plan: These past few days have been so intense that most crew members have not had the chance to finish the setup of their experiments. Now that our team is alone on Mars, everyone is excited to finally kick off their research without any interruptions.

Anomalies in work: Non-functional robotic telescope.

Weather: Sunny, but very high winds. Low 54F/High 71F.

Crew physical status: Everyone is feeling well, and our Crew Engineer is feeling much better than yesterday. She is nearly recovered!

EVA: We conducted an EVA to Marble Ritual in order to collect iron oxide samples.

Reports to be filed: Sol Summary, EVA request, EVA Report, HSO Report, Operations Report, Greenhab Report, Journalist Report, Astronomy Report, Pictures of the Day (with Photo of the Day)

Support Requested: We need new batteries for the gas detectors. We would also like to notify that all the stove knobs are broken and constitute a safety hazard (see HSO and Operations Report).

Copyright © The Mars Society. All rights reserved. | Main Site